OpenStudy (anonymous):

let v(vector space) be a collection of all 2*3 matrices with real entries such that v={[a11 a12 a13:a21 a22 a23]| a11 +a23 =1} determine whether the following vector spaces axioms hold on v .........still to post questions

OpenStudy (turingtest):

\[V=\left\{\left[\begin{matrix}a_{11}&a_{12}&a_{13}\\a_{21}&a_{22}&a_{23}\end{matrix}\right]|a_{11}+a_{23}=1\right\}\]correct?

OpenStudy (anonymous):

questions ..... (a)\[ni 0 vector in V\] such that \[alpha + 0 vector = 0 vector + alpha = alpha,forall alpha in V\]

OpenStudy (anonymous):

\[exists 0 vector in V\]

OpenStudy (turingtest):

Given\[V=\left\{\left[\begin{matrix}a_{11}&a_{12}&a_{13}\\a_{21}&a_{22}&a_{23}\end{matrix}\right]|a_{11}+a_{23}=1\right\}\]Find if\[\exists\vec0\in V:\vec\alpha+\vec0=\vec0+\vec\alpha=\vec\alpha,\forall\vec\alpha\in V\]right?

OpenStudy (turingtest):

I don't see how this could be true given that \(a_{11}+a_{23}=1\)

OpenStudy (anonymous):

oops...y is it not showing the equations??

OpenStudy (turingtest):

try reloading

OpenStudy (anonymous):

itz now ok....the plzvshow me da steps for saying itz not true

OpenStudy (turingtest):

in trying to get a zero vector I would have something like\[\vec0=\left[\begin{matrix}a_{11}&0&0\\0&0&1-a_{11}\end{matrix}\right]\]but clearly this will not give\[\vec0+\vec\alpha=\vec\alpha\]I don't know if that is enough proof for your prof but it convinces me lol

OpenStudy (anonymous):

i understand a lil bit ....but just dont get it how did u get that 0 (vector) matrix

OpenStudy (turingtest):

well I tried to make all elements zero, but given that\[a_{11}+a_{23}=1\]we have that\[a_{23}=1-a_{11}\]so all entries will be of the form\[\vec\alpha=\left[\begin{matrix}a_{11}&a_{12}&a_{13}\\a_{21}&a_{22}&1-a_{11}\end{matrix}\right]\]

OpenStudy (anonymous):

0⃗ +α⃗ =α⃗ from this equation i dont get it ..how will we represent vector 0...do we use other vector or just take 0 (vector ) as a 2*3 matrix containing all zeros

OpenStudy (turingtest):

normally yes, \[\vec0=\left[\begin{matrix}0&0&0\\0&0&0\end{matrix}\right]\]but since in this case \[a_{11}+a_{23}=1\implies a_{23}=1-a_{11}\]we know that all vectors will have the form\[\left[\begin{matrix}a_{11}&a_{12}&a_{13}\\a_{21}&a_{22}&1-a_{11}\end{matrix}\right]\]so the best we could do is\[\left[\begin{matrix}a_{11}&0&0\\0&0&1-a_{11}\end{matrix}\right]\]which doesn't help, so there exists no zero vector\[\vec0:\vec\alpha+\vec0=\vec\alpha\]

OpenStudy (anonymous):

i got prof ...thanks

OpenStudy (anonymous):

i had encountered a problem again when trying to implement 0+a=a ....it does not give me the solution dnt understand why

OpenStudy (turingtest):

that's because there is no zero vector in this space

OpenStudy (anonymous):

firstly should i take a +b=a then solve for b

OpenStudy (turingtest):

right, that should work

OpenStudy (turingtest):

\[\vec\alpha+\vec\beta=\vec\alpha\]\[\left[\begin{matrix}a_{11}&a_{12}&a_{13}\\a_{21}&a_{22}&1-a_{11}\end{matrix}\right]+\left[\begin{matrix}b_{11}&b_{12}&b_{13}\\b_{21}&b_{22}&1-b_{11}\end{matrix}\right]=\left[\begin{matrix}a_{11}&a_{12}&a_{13}\\a_{21}&a_{22}&1-a_{11}\end{matrix}\right]\]\[\left[\begin{matrix}b_{11}&b_{12}&b_{13}\\b_{21}&b_{22}&1-b_{11}\end{matrix}\right]=\left[\begin{matrix}0&0&0\\0&0&0\end{matrix}\right]\]this implies that\[b_{11}=1-b_{11}=0\]which is impossible, a contradiction. Hence there is no zero vector in this vector space.

OpenStudy (anonymous):

with the first 1 im ok....and the on same statement for verifying -a+a=0 ... does it holds for b23 =-a23

OpenStudy (turingtest):

if you want to verify -a+a=0 the you only need the fact that each element in a is the negative of each element in -a

OpenStudy (turingtest):

if you want you could write\[a_{23}=1-a_{11}\]\[-a_{23}=a_{11}-1\]\[a_{23}-a_{23}=1-a_{11}-(a_{11}-1)=0\huge~~~~~~~~~~~~~~~~~~~\checkmark\]but that is unnecessary.

OpenStudy (turingtest):

This also proves that the space is not closed under addition since we know that the zero vector is not in it. Hence this is not even a vector space.